OpenStudy (anonymous):

Explain, in complete sentences, how you would completely factor 20x2 – 28x – 48 by using the grouping method. How would you check your factors for accuracy?

OpenStudy (callisto):

First take out 4 as the common factor 20x^2 – 28x – 48 = 4 (5x^2 -7x -12) Then, consider (5x^2 -7x -48), break it into (5x^2 -12x +5x-12) Factor it by grouping method again (5x^2 -12x +5x-12) = 5x(x+1) -12(x+1) = (5x-12)(x+1) So, 20x^2 – 28x – 48 = 4 (5x^2 -7x -12) = 4(5x-12)(x+1) To check the accuracy, just expand it and see if you can get the terms given by the question Hope you understand it!

OpenStudy (anonymous):

Alright that makes sense, thankyou!

OpenStudy (callisto):

Welcome :)